OpenStudy (anonymous):

SOLVE FOR n n/30=12/72

pooja195 (pooja195):

We need to cross multiply so our set up would be : \[\huge~\rm~\frac{ n }{ 30 }=\frac{ 12 }{ 72 }\] Now we need to multiply so whats \(\huge\rm\color{red}{72 *~n=?}\) \(\huge\rm\color{red}{30*12=?}\)

OpenStudy (anonymous):

n=5

pooja195 (pooja195):

yes thats correct :)